OpenStudy (anonymous):

*medal to correct*A projectile is launched at a 40 degree angle . At the top of its motion, what is true?

OpenStudy (anonymous):

@ganeshie8 hello! could u answer/explain this short question? Would it be -9.8 acceleration and 0 Vy (velocity of y component) then why is it -9.8 acceleration??

OpenStudy (anonymous):

A. v = 0 m/s, a = 0 m/s^2 B. v_x = 0 m/s, a = 0 m/s^2 C. v_y = 0 m/s, a = -9.8 m/s^2 D. v = 0 m/s, a = -9.8 m/s^2

OpenStudy (anonymous):

Acceleration is always g and downward, if we ignore air drag.

ganeshie8 (ganeshie8):

g is never 0, when u r near by earth

ganeshie8 (ganeshie8):

also, v_x remains fixed during the flight as there is no acceleration in x direction